8. Nonconstant growth stock

As companies evolve, certain factors can drive sudden growth. This may lead to a period of nonconstant, or variable, growth. This would cause the expected growth rate to increase or decrease, thereby affecting the valuation model. For companies in such situations, you would refer to the variable, or nonconstant, growth model for the valuation of the company’s stock.

Consider the case of Portman Industries:

Portman Industries just paid a dividend of $2.40 per share. The company expects the coming year to be very profitable, and its dividend is expected to grow by 12.00% over the next year. After the next year, though, Portman’s dividend is expected to grow at a constant rate of 2.40% per year.

Assuming that the market is in equilibrium, use the information just given to complete the table.

Term

Value

Dividends one year from now (D₁)   
Horizon value (Pˆ1P̂1)   
Intrinsic value of Portman’s stock   

The risk-free rate (rRFrRF) is 3.00%, the market risk premium (RPMRPM) is 3.60%, and Portman’s beta is 1.10.

What is the expected dividend yield for Portman’s stock today?

4.77%

4.57%

4.46%

3.66%
Answer #1

a). D1 = D0 * (1 + g1) = $2.40 * (1 + 0.12) = $2.688

b). r = rRF + [Beta * RPM] = 3% + [1.10 * 3.60%] = 3% + 3.96% = 6.96%

D2 = D1 * (1 + gC) = $2.688 * (1 + 0.024) = $2.753

P1 = D2 / [r - gC] = $2.753 / [0.0696 - 0.024] = $2.753 / 0.0456 = $60.36

c). P0 = [D1 + P1] / [1 + r] = [$2.688 + $60.36] / [1 + 0.0696] = $63.05 / 1.0696 = $58.95

d). Expected Dividend Yield = D1 / P0 = $2.688 / $58.95 = 0.0457, or 4.57%

So, 2nd option is correct.